OpenStudy (anonymous):

If g and f(g(x)) are both onto, does it follow that f is onto? If g and f(g(x)) are both onto, does it follow that f is onto? @Mathematics

OpenStudy (jamesj):

Yes. In fact it is sufficient that (f o g)(x) = f(g(x)) is onto. Proof: I'll write these are functions U --> U, for some set U. As (f o g) is also onto, for all y in U, there is an x such that f(g(x)) = y. Now let w = g(x). Now as y was arbitrary, it follows that for every such y there is a number w such that f(w) = y.

OpenStudy (jamesj):

======== Example g(x) = sin x f(x) = tan(x.pi/2) g is only onto [-1,1], not the entire reals. But f(g(x)) = tan(sin x . pi/2) is onto the entire real line.